OpenStudy (anonymous):

ΔABC is similar to ΔPQR. AB¯¯¯¯¯ corresponds to PQ¯¯¯¯¯, and BC¯¯¯¯¯ corresponds to QR¯¯¯¯¯. If AB = 9, BC = 12, CA = 6, and PQ = 3, what are the lengths of QR¯¯¯¯¯ and RP¯¯¯¯¯?

OpenStudy (anonymous):

QR = 4, RP = 2 QR = 6, RP = 4.5 QR = 4.5, RP = 6 QR = 2, RP = 4

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

first option qr = 4 rp = 2

OpenStudy (anonymous):

well the ratio of similarity is AB/PQ = 9/3 = 3 ==> AC/RP=3 6/RP =3 HENCE RP=2CM BC/QR = 3 12/QR = 3 HENCE QR = 4 CM GET IT?? :P